2. luisa noticed that in 7 years she will be half her mothers age. if 3 yrs ago luisa was 1/3 her mothers age, how old is her mother now?
Answer by richwmiller(17219)

Website rules
Read them carefully. They all apply to you.
a) Limit 5 problems per day
b) No similar problems
c) Submit each problem one time separately and in only one category.
d= luisa m=mother
d+7=1/2(m+7)
d-3=1/3(m-3)
d = 13, m = 33